OpenStudy (anonymous):

Soybean meal is 12% protein ; cornmeal is 6% protein. How many pounds (lbs.) of each should be mixed together in order to get 240 lbs mixture that is 7% protein? How many lbs of cornmeal? How many lbs of soybean?

OpenStudy (anonymous):

something is missing....

OpenStudy (anonymous):

hold on let me re read it

OpenStudy (anonymous):

t it is not in my problem.

OpenStudy (anonymous):

let x for 12% =0.12 let y for 6% = 0.06 7%=0.07 x+y=240 x = 240-y 0.12x+ 0.06y= 0.07* 240 can you solve from now?

OpenStudy (anonymous):

17.8?

OpenStudy (anonymous):

subtitution x = 240-y 0.12(240-y)+ 0.06y= 0.07* 240 28.8 - 0.12y +0.06y = 16.8 -0.06y =16.7-28.8 y = -12/-0.06 y = 200

OpenStudy (anonymous):

x+y=240 subtitution y= 200 solve for x x+200=240 x = 240-200 x = 40 if I desever Medal than click good answer

OpenStudy (anonymous):

I don't get it lol

OpenStudy (anonymous):

How do I find out ho0w many pounds of each?

OpenStudy (anonymous):

cornmeal 40 lbs soybean 200lbs

OpenStudy (anonymous):

never mind I got it :D

OpenStudy (anonymous):

I told you let x for 12% =0.12 let y for 6% = 0.06 the answer x = 40 y =200 than : cornmeal 40 lbs soybean 200lbs u understand now? ty for medal

OpenStudy (anonymous):

yes I understand thank you :)

OpenStudy (anonymous):

uwc